What is the x intercept of the graph of the line that passes through the points (1,-15)and (2,-2)
Answer by nerdybill(7384)

What is the x intercept of the graph of the line that passes through the points (1,-15)and (2,-2)
.
slope = (y2-y1)/(x2-x1)
slope = (-2-(-15))/(2-1)
slope = (-2+15)/(1)
slope = 13
.
determine equation of line by plugging in info into "point-slope" form:
y - y1 = m(x -x1)
y - (-2) = 13(x - 2)
y + 2 = 13x - 26
y = 13x - 28
.
find x-intercept by setting y to zero:
y = 13x - 28
0 = 13x - 28
28 = 13x
28/13 = x
.
answer: (28/13, 0)